In a photovoltaic cell, the cathode is built from metallic Cu in a 1.0 M solution of Cu (NO3) 2, while the anode is an unknown metal introduced into a solution of its nitrate salt. If the total potential of the cell is 0.47 V and E0Cu = 0.34 V, what would be the standard potential of the unknown metal?
1
Expert's answer
2020-10-21T13:17:37-0400
"E^0(Me) = E^0(Cu\/Cu^{2+}) - E_{total}"

"E^0 = 0.34 - 0.47= -0.13V"